# DRPruning: Efficient Large Language Model Pruning through Distributionally Robust Optimization

### v1

@@ 5.1 Ablation Study
Furthermore, we analyze how our strategy dynamically adjusts training parameters, as illustrated in Figure 5. Both the 1.3B and 2.7B models show similar trends, increasing allocation to C4 and Wiki data and reduced allocation to CC, suggesting that our method provides reliable training expectations. This observation also underscores the limitations of static hyperparameter settings. The high potential of Wiki improvement leads to a significantly increased reference ratio, while in later training stages, slower loss convergence in the CC domain prompts a reduced weight.
Our method enables effective dynamic evaluation of reference loss.
To accelerate convergence, we select the minimum predicted value as the reference loss, which raises concerns about the inability to increase the reference loss when domain potential decreases. However, the two rightmost figures show that our predictions are conservative and consistently decrease during training. When potential declines, the rate of decrease slows, resulting in a relatively higher reference loss. This outcome stems from using the loss of a limited training step rather than the fully trained loss in scaling laws, resulting in cautious estimates.
@@
Table 4: Domain-level results under the benchmark we generated. The abbreviations CC, …, Stackexchange refer to the seven domains used for training in RedPajama.
5.2 Robustness across Different Domains
Table 5: Case For the sentence continuation task.
Our method outperforms the baselines consistently across the domains.
As shown in Table 4, our method consistently outperforms the Constant and Sheared Llama scheduling strategies, achieving better downstream performance in most domains. This demonstrates that our approach not only improves learning in hard domains but also preserves or enhances performance in others. Additionally, our benchmark results align well with the average performance across 15 tasks, performing slightly below the open-source Sheared Llama 2.7B model. This consistency confirms the quality and reliability of our benchmark and results.
@@ 5.3 Robustness under Distribution Shifts
To verify our method under larger distribution shifts, we conduct experiments under the multilingual setting.
Data setups.
We utilize the CulturaX dataset (Nguyen et al., 2024), a comprehensive multilingual resource that integrates mC4 and OSCAR, and has been rigorously cleaned and deduplicated. For our analysis, we focus on eight languages supported by Qwen2: English (EN), Russian (RU), Chinese (ZH), Japanese (JA), Arabic (AR), Turkish (TR), Korean (KO), and Thai (TH), which represent diverse language families.
@@
We use Qwen2 series models (Yang et al., 2024) as the base models and explore two approaches: (1) continued pretraining from Qwen2-1.5B, and (2) pruning Qwen2-7B and then continued pretraining. Due to grouped query attention differences, we keep the head dimension unchanged, resulting in a 1.8B target architecture. The reference loss is initialized with Qwen2-7B on the validation set. For the reference ratio, we follow Conneau et al. (2020), upsampling low-resource languages with a smoothing rate of 0.3. We select various downstream tasks and report average performance. Detailed model configurations and metrics are provided in Appendix A.2.
Our method demonstrates superior distribution robustness compared to Sheared Llama.
As shown in Table 6, our method outperforms Sheared Llama with an average gain of +1.50, including +3.48 in low-resource languages. It also surpasses the open-source Qwen2-1.5B, with an average gain of +2.98, and +6.14 in low-resource languages.
Our method dynamically allocates higher weights to low-resource languages.
As shown in Table 7, Sheared Llama focuses on hard-to-improve areas like English, where its loss remains above the reference loss, leading to a highly imbalanced data distribution. In contrast, our method dynamically identifies underperforming domains, increasing the reference loss for high-resource languages and lowering it for low-resource ones. This leads to more balanced data scheduling and better evaluation loss, further demonstrating the robustness of our approach in handling distribution shifts.
Continued pretraining from the pruned model underperforms from the pre-trained ones.
Continued pretraining from the pruned model result in a slight performance drop (average -0.48), despite the increased parameters (1.8B vs. 1.5B). This contrasts with Xia et al. (2024), where continued pretraining on a smaller model shows minimal gains. In our case, using different data and a stronger small model improves performance during continued pretraining, leading to different results.
@@ 6 Related Work
Pruning.
Unstructured pruning (Frankle and Carbin, 2019; Frantar and Alistarh, 2023; Sun et al., 2024) removes individual weights but offers limited speedup. This study focuses on structured pruning (Han et al., 2015; Wen et al., 2016), which removes entire structural components, such as neurons or layers, making it more effective for simplifying models and improving efficiency. In task-specific models, extensive pruning can retain performance (Liu et al., 2017; Wang et al., 2020a; Lagunas et al., 2021; Xia et al., 2022; Kurtic et al., 2023). However, for large language models (LLMs), as the ratio of training data to model size increases (Hoffmann et al., 2022b; Touvron et al., 2023a, b), fewer redundant parameters remain, leading to significant performance degradation after pruning.
To counter this, performance recovery techniques like continued pretraining are essential (Ma et al., 2023; Zhang et al., 2024b). However, continued pretraining of pruned models reduces loss at different rates across domains compared to training from scratch, resulting in less efficient data utilization (Xia et al., 2024). To address this, we employ a distributionally robust pruning method that dynamically adjusts the data distribution during training, ensuring balanced performance across domains.
@@
Overparameterized neural networks often perform well on i.i.d. test sets but struggle with underrepresented data groups (Hovy and Søgaard, 2015; Blodgett et al., 2016; Tatman, 2017; Duchi et al., 2020). To address this, unlike empirical risk minimization (ERM), which minimizes expected loss for a fixed distribution, MultiDDS (Wang et al., 2020b) uses a gradient-based meta-learning method to optimize the sampling distribution. However, MultiDDS increases computational and memory costs. In contrast, DRO (Delage and Ye, 2010; Ben-Tal et al., 2012; Bertsimas et al., 2014) improves performance without additional complexity (Hashimoto et al., 2018; Oren et al., 2019; Levy et al., 2020).
@@
DRO shows strong performance but relies on two main hyperparameters. The first is the reference loss, usually set by training an additional baseline model (Zhou et al., 2021; Jiao et al., 2022), though this is expensive for LLMs. Xia et al. (2024) attempt to reduce costs by using scaling laws to predict the pruned model’s performance, but our experiments show this method fails under severe distribution shifts during continued pretraining. The second hyperparameter is the reference data ratio, often determined through temperature-based sampling (Arivazhagan et al., 2019; Conneau et al., 2020) or manually (Touvron et al., 2023a; Parmar et al., 2024). However, fixed ratios can hinder model convergence in challenging distributions. We propose a dynamic method that shifts weight toward higher-loss domains, enhancing distribution robustness and improving downstream performance.
@@ 7 Conclusion
We present DRPruning, a distributionally robust pruning method that addresses uneven performance degradation across domains during structured pruning. By utilizing and further improving distributionally robust optimization (DRO), our method focuses more on domains with poorer performance, accelerating performance recovery. It outperforms existing models and training methods in both monolingual and multilingual settings, achieving lower perplexity, higher task accuracy, and better instruction tuning. Further analysis demonstrates the robustness of our method against various domains and distribution shifts. Additionally, the dynamic adjustment of reference loss and data ratios exhibits broad applicability, with strong potential to support balanced training across diverse tasks.
@@
Due to computational constraints, we are unable to explore pruning to larger models, i.e., employing smaller pruning ratios. Retaining a larger proportion of the model’s parameters may lead to different outcomes in some experiments. For example, it remains to be investigated whether pruning larger models provides benefits, and whether it is better to continue pretraining from a pruned model or from a smaller, fully pre-trained model.
@@
Xia et al. (2024) point out that pruned models exhibit higher training ceilings. Although good performance can be achieved with tens of billions of training samples, this study does not investigate whether training the models to full convergence using hundreds or thousands of billions of samples would yield better results than continuing pretraining from existing pre-trained models under similar settings.
@@
We have validated our method’s effectiveness in the pruning phase, the pruning recovery phase, and the continued pretraining phase. However, our method is expected to be applicable in broader contexts, such as pretraining from scratch and cross-domain instruction tuning. Broader validation would further demonstrate the superiority of our approach.
@@ A.1 Main Experiment
For the DRO setup, we follow Zhou et al. (2021) closely. The constraint size $\rho$ for the chi-square ball is set to {0.05, 0.1, 0.2}. Preliminary experiments show that $\rho=0.1$ yields the best results, so we use this value in all experiments. Following their setup, we truncate the dynamic data ratio to prevent it from dropping below the minimum reference data ratio, which further ensures balanced domain training. We compute historical loss values using an exponential moving average, with the hyperparameter $\lambda$ set to 0.1, which is also used for updating the reference data ratio. Besides, for the prediction of the reference loss, we maintain an average loss below $3\times 10^{-5}$, demonstrating the effectiveness of our method.
@@ B.1 Analysis of Mask Similarity
Table 11: Mask similarity mean values and the Standard Error of the Mean (SEM) under different data scheduling strategies and random seeds. “Same” indicates using identical data scheduling but different random seeds, while “Different” indicates using different data scheduling strategies.
@@
Table 12: The performance of pruning Llama2-7B and 13B models down to 1.3B and 2.7B parameters. “Pruning” refers to using the pruned model without continued pretraining, while “Continue Pretraining” means using the model after continued pretraining. “From” and “To” indicate the size of the source and target model, respectively.
@@
We analyze mask similarity across training sessions with different random seeds under identical and distinct data scheduling strategies. The results, shown in Table 11, present mean values and Standard Error of the Mean (SEM). The trend across pruning dimensions aligns with the previous findings: similarity is higher at the coarse-grained layer level and lower at the finer-grained intermediate level. Besides, comparisons across three pairs under identical settings and twelve pairs under different settings show consistently low similarity. However, no significant difference in perplexity is observed, suggesting that the model’s interchangeable parameters allow similar outcomes despite different pruning decisions. This randomness obscures variations caused by differing data distributions.
@@ B.2 Pruning from Larger LLMs
We investigate whether pruning from larger language models (LLMs) with a higher pruning ratio provides additional benefits. Experiments are conducted in the monolingual setting, consistent with the main text, to compare the effects of pruning from Llama2-7B and Llama2-13B.

### v2

@@ 5.1 Ablation Study
Table 5: Domain-level results under the benchmark we generated. The abbreviations of tasks refer to the evaluation of seven domains used for training in RedPajama.
@@
5.2 Robustness across Different Domains
DRPruning outperforms the baselines consistently across the domains.
As shown in Table 5, our method consistently outperforms the Constant and Sheared Llama scheduling strategies, achieving better downstream performance in most domains. This demonstrates that our approach not only improves learning in hard domains but also preserves or enhances performance in others. Additionally, our benchmark results align well with the average performance across 15 tasks, performing slightly below the open-source Sheared Llama 2.7B model. This consistency confirms the quality and reliability of our benchmark and results. Furthermore, we demonstrate that DRPruning benefits from finer domain segmentatio in Appendix B.5.
Figure 6: Variation of data ratio, reference data ratio, and reference loss during continued pretraining. The top three plots show results for pruning to 1.3B parameters, while the bottom three are for pruning to 2.7B parameters. Dashed lines in data ratio plots represent the initial reference data ratio for each domain. Reference loss plots display the difference from the initial value, with the dashed line at $y=0$ indicating no change from the initial ones.
@@ 5.3 Robustness under Distribution Shifts
To verify our method under larger distribution shifts, we conduct experiments under the multilingual setting. To ensure a fair comparison, we reproduce DRPruning and ReSheared methods under the same experimental setup, detailed as follows:
@@
We use Qwen2 series models (Yang et al., 2024), which demonstrate superior multilingual performance, as the base models and explore two approaches: (1) continued pretraining from Qwen2-1.5B, and (2) pruning Qwen2-7B and then continued pretraining. Due to grouped query attention differences, we keep the head dimension unchanged, resulting in a 1.8B target architecture. We use the CulturaX dataset (Nguyen et al., 2024) and select eight languages covered by Qwen2. The reference loss is initialized with Qwen2-7B on the validation set. For the reference ratio, we follow Conneau et al. (2020), upsampling low-resource languages with a smoothing rate of 0.3. We select various downstream tasks and report average performance. Detailed model configurations and metrics are provided in Appendix A.2.
Our method demonstrates superior distribution robustness compared to ReSheared.
As shown in Table 6, our method outperforms ReSheared with an average gain of +1.50, and Qwen2-1.5B with +2.98. To analyze the source of these gains, as shown in Table 7, Sheared Llama focuses on hard-to-improve areas like English, where its loss remains above the reference loss, leading to a highly imbalanced data distribution. In contrast, our method dynamically identifies underperforming domains, increasing the reference loss for high-resource languages and lowering it for low-resource ones. This leads to more balanced data scheduling and better evaluation loss, further demonstrating the robustness of our approach in handling distribution shifts.
Continued pretraining from the pruned model underperforms from the pretrained ones.
Continued pretraining from the pruned model result in a slight performance drop (average -0.48), despite the increased parameters (1.8B vs. 1.5B). This contrasts with Sheared Llama, where continued pretraining on a smaller model shows minimal gains. In our case, using different data and a stronger small model improves performance during continued pretraining, leading to different results. Furthermore, DRPruning demonstrates efficacy not only under continued pretraining but also under instruction tuning, detailed in Appendix B.6.
5.4 Analysis of Hyperparameter Adjustment
We analyze how our strategy adjusts training parameters, including the data ratio, reference data ratio, and reference loss, in the main experiment. The results are shown in Figure 6.
High reliability of our approach.
The trends for the 1.3B and 2.7B targets are similar, with increased allocation to C4 and Wiki and reduced allocation to CC. This highlights limitations in the current hyperparameter settings while confirming the reliability of our dynamic scheduling. In the later training stages, the CC domain exhibits lower potential with slower loss convergence, prompting our strategy to reduce its weight and increase the weight for C4. Wiki data consistently shows higher potential, leading to a significantly higher reference data ratio and the largest reference loss reduction.
Effective real-time evaluation of reference loss.
To accelerate convergence, we select the minimum predicted value, which raises concerns about the inability to increase the reference loss when domain potential decreases. However, the two rightmost figures show that our predictions are conservative and decrease gradually during training. When potential declines, the rate of decrease slows, resulting in a relatively higher reference loss. This favorable outcome arises because we use loss from a limited training duration instead of the fully trained loss used in scaling laws, leading to more cautious estimates. Further, the similar trends between the 1.3B and 2.7B models indicate that our method provides reasonable training expectations, potentially supporting model training across broader ranges.
@@ 6 Related Work
LLM Pruning.
Unstructured pruning (Frankle and Carbin, 2019; Frantar and Alistarh, 2023; Sun et al., 2024) removes individual weights but offers limited speedup. This study focuses on structured pruning (Han et al., 2015; Wen et al., 2016), which removes entire structural components, making it more effective for improving efficiency Liu et al. (2024); Rao et al. (2023). In task-specific models, extensive pruning can retain performance (Liu et al., 2017; Wang et al., 2020a; Lagunas et al., 2021; Xia et al., 2022; Kurtic et al., 2023). However, for LLMs, as training data increases (Hoffmann et al., 2022b), fewer redundant parameters remain, leading to significant performance degradation after pruning. To counter this, performance recovery techniques like continued pretraining are essential (Ma et al., 2023; Zhang et al., 2024b). However, continued pretraining of pruned models reduces loss at different rates across domains, resulting in less efficient data utilization (Xia et al., 2024). To address this, DRPruning dynamically adjusts the data distribution during training, ensuring balanced performance across domains.
@@
Overparameterized neural networks excel on i.i.d. test sets but struggle with underrepresented data groups (Hovy and Søgaard, 2015; Blodgett et al., 2016; Tatman, 2017). Unlike empirical risk minimization, which minimizes expected loss for a fixed distribution, MultiDDS (Wang et al., 2020c) optimizes the sampling distribution via gradient-based meta-learning but incurs higher computational and memory costs. In contrast, DRO (Delage and Ye, 2010; Ben-Tal et al., 2012; Bertsimas et al., 2014) improves performance without additional complexity (Hashimoto et al., 2018).
@@
DRO shows strong performance but relies on two main hyperparameters. The first is the reference loss, usually set by training an additional baseline model (Zhou et al., 2021; Jiao et al., 2022), though this is expensive for LLMs. Sheared Llama uses scaling laws of model size to predict the pruned model’s performance. The second hyperparameter is the reference data ratio, often determined through temperature-based sampling (Arivazhagan et al., 2019; Conneau et al., 2020) or manually (Touvron et al., 2023a; Parmar et al., 2024). However, fixed ratios can hinder model convergence in challenging distributions. DRPruning shifts weight toward higher-loss domains, enhancing distribution robustness and improving downstream performance.
@@ 7 Conclusion
This paper presents DRPruning, a distributionally robust pruning method that addresses uneven performance degradation across domains during structured pruning. By utilizing and further improving distributionally robust optimization (DRO), our pruning method focuses more on domains with poorer performance, significantly accelerating performance recovery. It outperforms existing models and data scheduling methods in both monolingual and multilingual settings, achieving lower perplexity, higher task accuracy, and better instruction tuning outcomes. Further analysis demonstrates the robustness of our method against various domains and distribution shifts. Additionally, the dynamic adjustment of reference loss and data ratios exhibits broad applicability, with strong potential to support balanced training across diverse tasks.
@@
Due to computational constraints, we are unable to explore pruning to larger models, i.e., employing smaller pruning ratios. Retaining a larger proportion of the model’s parameters may lead to different outcomes in some experiments. For example, it remains to be investigated whether pruning larger models provides benefits, and whether it is better to continue pretraining from a pruned model or from a smaller, fully pretrained model.
@@
Xia et al. (2024) point out that pruned models exhibit higher training ceilings. Although good performance can be achieved with tens of billions of training samples, this study does not investigate whether training the models to full convergence using hundreds or thousands of billions of samples would yield better results than continuing pretraining from existing pretrained models under similar settings.
@@
We have validated our method’s effectiveness in the pruning phase, the pruning recovery phase, the continued pretraining phase, and the instruction tuning phase in Appendix B.6. Our method is expected to be applicable in broader contexts, such as pretraining from scratch and cross-domain post-training. Broader validation would further demonstrate the superiority of our approach.
@@
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62206076), Guangdong S&T Program (Grant No. 2024B0101050003), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515011491), and Shenzhen Science and Technology Program (Grant Nos. ZDSYS20230626091203008, KJZD20231023094700001, KQTD2024072910215406). We would like to thank the anonymous reviewers and meta-reviewer for their insightful suggestions.
@@ A.1 Main Experiment
We follow Sheared Llama to update the data ratio every 50 steps during pruning and every 400 steps during continued pretraining. For the DRO setup, we follow Zhou et al. (2021) closely. The constraint size $\rho$ for the chi-square ball is set to {0.05, 0.1, 0.2}. Preliminary experiments show that $\rho=0.1$ yields the best results, so we use this value in all experiments. Following their setup, we truncate the dynamic data ratio to prevent it from dropping below the minimum reference data ratio, which further ensures balanced domain training. We compute historical loss values using an exponential moving average, with the hyperparameter $\lambda$ set to 0.1, which is also used for updating the reference data ratio. Besides, for the prediction of the reference loss, we maintain an average loss below $3\times 10^{-5}$, demonstrating the effectiveness of our method.
@@ A.2 Multilingual Experiment
Data.
We use the CulturaX dataset (Nguyen et al., 2024), a large multilingual resource with 6.3 trillion tokens across 167 languages, integrating mC4 and OSCAR, and meticulously cleaned and deduplicated. We select eight languages covered by Qwen2: English (EN), Russian (RU), Chinese (ZH), Japanese (JA), Arabic (AR), Turkish (TR), Korean (KO), and Thai (TH), representing diverse language families.
@@ B.1 Analysis of Mask Similarity
Table 11: Mask similarity mean values and the standard error of the mean under different data scheduling strategies and random seeds. “Same” indicates using identical data scheduling but different random seeds, while “Different” indicates using different data scheduling strategies.
@@
Table 12: The performance of pruning Llama2-7B and 13B models down to 1.3B and 2.7B parameters. “Pruning” refers to using the pruned model without continued pretraining, while “Continue Pretraining” means using the model after continued pretraining. “From” and “To” indicate the size of the source and target model, respectively. Bold indicates superior performance when pruning from 7B or 13B.
@@
We analyze mask similarity across training sessions with different random seeds under identical and distinct data scheduling strategies. The results, shown in Table 11, present mean values and standard error of the mean. The trend across pruning dimensions aligns with the previous findings: similarity is higher at the coarse-grained layer level and lower at the finer-grained intermediate level. Besides, comparisons across three pairs under identical settings and twelve pairs under different settings show consistently low similarity. However, no significant difference in perplexity is observed, suggesting that the model’s interchangeable parameters allow similar outcomes despite different pruning decisions. This randomness obscures variations caused by differing data distributions.
@@ B.2 Pruning from Larger LLMs
We investigate whether pruning from LLMs with a higher pruning ratio provides additional benefits. Experiments are conducted in the monolingual setting, consistent with the main text, to compare the effects of pruning from Llama2-7B and Llama2-13B.